2. Classify the following as strong or weak acids/bases. Write balances equations that describe the dissociation of the compounds in water. 

a. HF b. NH3 c. NaOH d. C6H5NH2 e. HNO2 f. CH3COOH g. HCI h. H2CO3 


3. Please the species in each of the following groups in order of increasing acid strength. 

a. HIO3, HCIO3, HBrO3 b. HF, HI, HCI c. HOCI, HOI, HOF e. H2O, H2S, H2Se f. HBO, HBrO2, HBrO4, HBrO3
Answer #1

a) HF - Weak acid

HF+ H2O = H3O+ + F-

b) NH3 = Strong base

NH3 + H2O = NH4+ + OH-

c ) NaOH = Strong acid

NaOH + H2O = Na+ (aq) + OH- (aq)

d) C6H5NH2 = Weak base

C6H2NH2 + H2O = C6H5NH3+ + H3O+

e ) HNO2 = Weak acid

HNO2 + H2O = H3O+ + NO2-

f) CH3COOH = Weak acid

CH3COOH + H2O = CH3COO- + H3O+

g) HCl = Strong acid

HCl + H2O = Cl- +H3O+

h) H2CO3 = Weak acid

H2CO3 + H2O = H3O+ + HCO3- = 2H3O+ + CO3-

3. a) HIO3, HClO3, HBrO3 : Acidity of these acids depend upon strength electronegativity of halogen. Cl being the highest electronegative atom, it pulls the electron of OH bond towards itself. So, release more proton. So, HClO3 is most acidic among these acids. Electronegativity order of halogen is Cl > Br>I. So acidity order HClO3>HBrO3>HIO3.

b) Acidity of halogen acid depends on bond strength of H-X. With increasing size of halogen, H-X bond become weaker. Bond strength order of halogen is HF> HCl> HI. Hence acidity order HI>HCl>HF.

c) HOCl, HOI, HOF : Acidity of Oxy acid depends on polarity of OH bond. Polarity of OH bond depends on electronegativity of attached halogen. More electronegative halogen make the acid more acidic. So acidity order : HOF> HOCl> HOI

e) H2O , H2S, H2Se : Here acidity depends on X-H bond (X= O,S, Se). As the size of S increases, X-H bond become weaker due to poorer overlap. So acidity order H2Se> H2S> H2O

f) HBrO, HBrO2,  HBrO4, HBrO3 : For oxyacid with same central atom, acidity increases with increase in no of oxygen atom, So acidity order HBrO4> HBrO3> HBrO2> HBrO.